OpenStudy (anonymous):

the area of a circle is 16 pi. what is the circumference?

OpenStudy (anonymous):

8pi

OpenStudy (anonymous):

Because A=16pi , A=r^2 . r=\sqrt{16}=4. C=2*4pi=8pi.